Two reservoirs are connected by a 5000 m long and 1.5 m diameter C.I pipe with f = 0.01 . The difference in water levels in two reservoir is 25 m. Determine the increase in discharge if the C.I pipe is replaced with a smooth steel pipe of same size with f = 0.005.


1
Expert's answer
2021-06-17T07:37:38-0400

"H_L=\dfrac{32f_1LQ_1²}{π²gd_1
⁵}"


"Q_1= \sqrt{\dfrac{Hπ²gd_1
⁵}{32f_1L}}"


"Q_1 = \sqrt{\dfrac{25π²×9.81×1.5⁵}{32×0.01×500}}= 10.72m³/s"


"Q_2= \sqrt{\dfrac{Hπ²gd_2
⁵}{32f_2L}}"


"\sqrt{\dfrac{25π²×9.81×1.5⁵}{32×0.005×500}} = 15.16m³/s"


"Q= Q_1+ Q_2=15.16m\u00b3\/s- 10.72m\u00b3\/s= 4.44m\u00b3\/s"